# Context-Aware Deep Learning for Robust Channel Extrapolation in Fluid Antenna Systems

### v1

@@ I Introduction
Since the introduction of FAS by Wong et al. in [8, 9], there have been lots of efforts in understanding the achievable performance of FAS [10, 11]. The use of FAS has also been considered for secrecy communications [12], reconfigurable intelligent surface (RIS) systems [13], index modulation [14, 15], wideband communication [16], and integrated sensing and communication (ISAC) [17]. Capacity maximization for FAS channels has also been studied using statistical channel state information (CSI) [18] and instantaneous CSI [19]. The diversity-multiplexing trade-off for multiple-input multiple-output (MIMO) FAS channels was analyzed in [20].
Nonetheless, the promising performance of FAS relies on the availability of CSI and the high-resolution nature of FAS also leads to substantial overhead in acquiring CSI, including pilot overhead, feedback latency, and energy consumption. In recent years, several studies explored the integration of physical modeling and data-driven approaches for CSI prediction [21, 22]. For instance, [23] proposed an inference algorithm to enable high-accuracy extrapolation in a massive FAS using only a small number of observed ports. Additionally, in [24], an extrapolation scheme based on masked language models, which reconstructs the full channel from a limited number of observed port CSI, was proposed. Going one step further, the approach in [25] integrates the diffusion mechanism of graph neural networks with an asymmetric masked autoencoder architecture, significantly improving both extrapolation accuracy and flexibility in high-resolution FAS settings. Prior to these, there were attempts that exploited channel sparsity [26, 27] and used Bayesian approaches [28, 29]. CSI prediction for multiuser FAS using deep learning was also reported in [30]. Overall, however, reliable CSI acquisition seems only possible at high signal-to-noise ratio (SNR) and it is important to study enhancement techniques that work well in noisy situations.
@@ 2 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $2$ LeakyReLU
3 CAB $32~{}\times~{}6$ $16~{}\times~{}8$ $128$ $3$ $1$ LeakyReLU

### v2

@@ I Introduction
Since the introduction of FAS by Wong et al. in [8, 9], there have been lots of efforts in understanding the achievable performance of FAS [10, 11]. The use of FAS has also been considered for reconfigurable intelligent surface (RIS) systems [12], index modulation [13, 14], wideband communication [15], and integrated sensing and communication (ISAC) [16]. Capacity maximization for FAS channels has also been studied using statistical channel state information (CSI) [17] and instantaneous CSI [18]. The diversity-multiplexing trade-off for multiple-input multiple-output (MIMO) FAS channels was analyzed in [19].
Nonetheless, the promising performance of FAS relies on the availability of CSI and the high-resolution nature of FAS also leads to substantial overhead in acquiring CSI, including pilot overhead, feedback latency, and energy consumption[20]. In recent years, several studies explored the integration of physical modeling and data-driven approaches for CSI prediction [21, 22]. For instance, [23] proposed an inference algorithm to enable high-accuracy extrapolation in a massive FAS using only a small number of observed ports. Additionally, in [24], an extrapolation scheme based on masked language models, which reconstructs the full channel from a limited number of observed port CSI, was proposed. Going one step further, the approach in [25] integrates the diffusion mechanism of graph neural networks with an asymmetric masked autoencoder architecture, significantly improving both extrapolation accuracy and flexibility in high-resolution FAS settings. Prior to these, there were attempts that exploited channel sparsity [26] and used Bayesian approaches [27]. CSI prediction for multiuser FAS using deep learning was also reported in [28]. Overall, however, reliable CSI acquisition seems only possible at high signal-to-noise ratio (SNR) and it is important to study enhancement techniques that work well in noisy situations.
@@ 2 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $2$ LeakyReLU
3 CAB $32~{}\times~{}16$ $16~{}\times~{}8$ $128$ $3$ $1$ LeakyReLU